If Coa x=sin y & cot (x-40^0)=tan (50^0-y). Find the value of x & y.

Grade:10

1 Answers

Arun
6 years ago
Dear Ratan
 
cos x = cos(90-y)
Hence x+y = 90° or we can say x and y are complementary angle of each other.
Now
 
cot(90-y-40) = cot(90-(50-y))
Now on comparison
90-y-40 = 90-50+y
2y = 10
Hence y = 5° 
& x = 85°
 
Regards
Arun (askIITians forum expert)